OpenStudy (anonymous):

(2x+5)^-(x-3)^, I see as a difference of squares, but how do I know what to factor?

zepdrix (zepdrix):

(2x+5)^2 - (x-3)^2 Is this what the problem is suppose to look like? :)

zepdrix (zepdrix):

\[a^2 - b^2 = (a-b)(a+b)\]\[(2x+5)=a\] Does that help you get started? :D I know it can be alil confusing with all the brackets!

OpenStudy (anonymous):

Yes, thank you very much